Calculate the molality and molarity of commercial ammonium hydroxide given the following data:NH3: weight%= 25.00 ; specific gravity=0.96g/ml
Answer #1

25% NH3 means 25gm of NH3 present in 100gm of water

molality    = weight of solute*1000/gram molar mass * weight of solvent in gm

              = 25*1000/35*100

              = 250/35 = 7.14 m

Molarity   = percentage *specific gravity*10/gram molar mass

            = 25*0.96*10/35 = 6.65 M